# Dynamic borrowing from historical controls via the synthetic prior with covariates in randomized clinical trials

### v1

@@ 3 Adaptive Design Based on Posterior Inference
We propose a two-stage adaptive design, largely following Schmidli et al. (2014), to reduce the control group size to the extent that reliable information can be gained through historical borrowing. Intuitively, in Stage 1 the new trial enrolls a fixed and prespecified number of control group patients. Then there is an interim check to determine the number of control patients to enroll in Stage 2. The interim check measures the degree of “compatibility” between the historical and new trial (Stage 1) data in terms of an effective sample size of control patients gained by borrowing from the historical data. If the new trial’s stage one control data is deemed less compatible with the historical data, little borrowing will happen and the Stage 2 size will not be reduced much or at all. This limits the impact of prior assumptions about the relevance of the historical data on the overall trial size. Because we model the treatment group data independently from the control group data (having no shared parameters), the treatment group patients may be enrolled without reference to the control patients or the stages of the adaptive design (i.e. by simply changing the randomization probability after the interim check).
@@
    Collect data on $n^{c}_{1}$ control patients.
  • Interim:
    Calculate the prior effective sample size $n^{c}_{eff}$ of the SPx model for $\psi_{H+1}$ given the Stage 1 data.
@@
    Collect data on $n^{c}_{2}$ more control patients, where $n^{c}_{2}=n^{c}_{max}-n^{c}_{eff}-n^{c}_{1}$ but truncated to be at least $p_{min}n^{c}_{max}$ but no more than $p_{max}n^{c}_{max}$ for some $p_{min}\in[0,1]$ and $p_{max}\in[1,\infty)$ (e.g. $p_{min}=0.75$ and $p_{max}=1.25$).
The second stage is designed so that the total sample size of the control group is not intolerably lower or higher than the target size $n^{c}_{max}$.
Crucial to this design is the prior effective sample size of SPx at interim, $n^{c}_{eff}$, which is intended to assess how much information the historical data contribute about the new trial’s control rate and how many fewer patients the new control group needs in exchange for this additional information (Hobbs et al. 2013, Schmidli et al. 2014). We use a simple moment matching definition of effective sample size (Weber 2020), which finds the Beta distribution with the same mean and variance as the SPx posterior, takes the effective sample size of this Beta distribution (the sum of its parameters; Morita et al. 2008), and subtracts the current sample size at interim, $n^{c}_{1}$. We note that it is technically possible that $n^{c}_{eff}-n^{c}_{1}<0$, indicating that using the historical data reduce certainty about the new trial’s control rate and the Stage 2 sample size should be slightly increased to compensate. In truth, defining an effective sample size that has desirable properties for complex non-conjugate models such as SPx is an open area of research (Morita et al. 2008, Morita et al. 2012, Neuenschwander et al. 2020). However, the sample size thresholds used in Stage 2 of the design somewhat limit the impact of the choice of definition. The simple definition we use has the benefit of producing more conservative (smaller) effective sample sizes for SPx than other definitions, which sometimes produce implausibly large effective sample sizes for the SPx model. Regardless of the specific definition used, by measuring the effective sample size during, and not before, the trial we can measure the extent to which the new data diverge from the historical data, potentially safeguarding against inappropriate borrowing from irrelevant historical data.
@@
$P(\delta>0|D)$ $\geq 1-q_{positive}$ (3)
for the treatment effect $\delta:=\psi_{trt}-\psi_{H+1}$ where $D$ is all of the historical and new trial data and $q_{positive}\in(0,1)$ is the posterior Type I error threshold. Alternatively, to detect clinically significant effects, we may use the rule
$P(\delta>\delta_{0}|D)$ $\geq 1-q_{clinical}$ (4)
@@ 4 Simulation Study
We compared several methods including SPx, RMAP (Schmidli et al. 2014), and an independent model (Ind.) with no historical borrowing (i.e. $\theta_{H+1}\sim Logistic(0,1)$). For SPx and RMAP we included both versions where the new trial’s control group size was fixed and where the two-stage adaptive design described in Section 3 was used (with $n^{c}_{1}=n^{c}_{max}/2$, $p_{max}=1.25$, and $p_{min}=0.75$). The two versions gave a contrast on the potential gain in the adaptive design in reducing the control group sample size. The value $n^{c}_{max}$ is fixed at 150 or 80, which are the maximum control sample size. For SPx our standard implementation used only 2 of 6 covariates associated with response rates to mimic imperfect knowledge or data collection, but we also include cases where SPx uses all 6 covariates to show the effectiveness of the method when the regression is strong and not misspecified. For RMAP we used a 50-50 mixture between the borrowing (MAP) component and the non-borrowing component, since this specification performed well in Schmidli et al. (2014). The 90-10 mixture they consider was too aggressive and extremely biased in many of our scenarios.
@@ 4.1 Estimation Accuracy for the New Trial’s Control Rate
RMSE 0.018 0.018 0.026 0.028 0.029 0.022 0.022 0.031 0.033 0.037
Coverage 98.8 98.9 95.0 96.4 93.5 99.3 99.2 97.6 97.8 96.3
Width 0.086 0.094 0.103 0.112 0.110 0.117 0.130 0.138 0.152 0.154
@@
Table 1: Control group size and Frequentist estimation accuracy for the new trial’s control response rate, averaged over 1,000 simulated trials. Metrics are defined as follows: size is the mean control group size; RMSE is the root mean square error of the posterior mean of the control group rate; coverage is the proportion of trials in which the 95% quantile-based posterior credible interval for the control group rate contains the true rate; width is the mean width of these credible intervals. Note that the credible intervals are Bayesian and not designed or calibrated to give exactly 95% Frequentist coverage.
@@ 4.2 Power and Type I Error for the New Trial’s Treatment Effect
Table 2 examines the statistical power of the adaptive designs assuming 2 and 6 covariates were included in SPx, respectively. Type I error is relatively well controlled by all methods. Of course, a drug developer or regulator requiring that Frequentist Type I error be more strictly controlled can calibrate the model or decision rule by simulation under the specific scenarios they are concerned about. This is a reality of all Bayesian trial methods and some Frequentist ones as well (e.g. Lewis et al. 2007), a point we revisit in the Discussion.
@@
0 $P(\delta>0)$ 5.5 5.2 6.6 5.9 5.0 6.2 6.3 8.2 8.8 5.1
0.3 $P(\delta>0.2)$ 69.8 70.2 70.1 71.4 69.5 45.2 45.7 47.8 50.1 43.0
Table 2: Frequentist Type I error and power for the new trial’s treatment effect, averaged over 1,000 simulated trials. For convenience, the mean control group size is repeated from Table 1. Type I error is shown in the $P(\delta>0)$ rows, where each model declares a positive treatment effect when its posterior probability that $\delta>0$ exceeds 95%; this row shows the % of simulated trials where $\delta=0$ in which each method instead declares $\delta>0$. Power is shown in the $P(\delta>0.2)$ rows, where each model declares a clinically significant treatment effect when its posterior probability that $\delta>0.2$ exceeds 95%; this row shows the % of simulated trials where $\delta=0.3$ in which each method declares $\delta>0.2$.
@@ 4.3 SPx’s Adaptive Weighting of Borrowing Mechanisms
To illustrate how SPx automatically adjusts the type of borrowing it performs depending on the historical and new trial data, Figure 1 plots the simulation distribution of SPx’s posterior submodel weights $p^{*}_{k}$ (see equation (2)) in Scenarios 1 through 4. When the historical data are entirely misleading (Scenario 4) SPx strongly favors its no-borrowing submodel. Otherwise SPx puts most posterior mass on its two borrowing submodels, appropriately favoring the regression submodel over the historical one when the historical rates are misleading but covariates are useful (Scenario 2). In the reversed setting where the historical rates are not misleading but covariates are not predictive (Scenario 3) SPx still gives moderate weight to the regression model because its inclusion of covariates adds some noise but not substantial bias. This can be verified from small values of the estimated regression coefficients (results not shown).
@@
5 Case Study
@@
We discuss the application of SPx to the development of novel treatments for rheumatoid arthritis (RA). RA is an auto-immune disease that affects more than 1.3 million patients in the United States (Hunter et al. 2017), and it is predicted that its global burden will increase through 2030 (Cai et al. 2023). The symptoms of this disease include pain and swelling joints in hands and feet as well as morning stiffness lasting longer than 30 minutes. Although there is no cure for RA, several treatments are available to slow down the disease progression and alleviate symptoms. Adalimumab is a well-established, standard biologic therapy in RA and has been approved for 20 years. As the patent of adalimumab is expiring globally and becoming a standard of care, novel therapies for RA must demonstrate superior treatment effects over adalimumab. We apply our SPx methodology in adalimumab trials conducted in the past two decades to explore possibilities for the design and analysis of a future trial in RA. Since adalimumab’s own development was extensive, there are many past trials including an adalimumab arm that could be used as a rich source of historical data to potentially accelerate trials of novel RA drugs. In this process, SPx uses trial-level covariate information and a combination of historical borrowing mechanisms to guide the new trial.
@@
ALTARA Kennedy et al. (2014) MTX 48.8 43 39.5
ARMADA Weinblatt et al. (2003) MTX 56.0 62 21.0
DE019 Keystone et al. (2004) MTX 56.1 200 24.0
IM133-001 Weinblatt et al. (2015) MTX 51.4 61 39.3
ORAL-Standard van Vollenhoven et al. (2012) MTX 53.7 106 26.4
RA-BEAM Taylor et al. (2017) MTX 53.0 488 40.2
STAR Furst et al. (2003) MTX 55.8 315 29.5
A3921035 Fleischmann et al. (2012) none 53.0 59 22.0
CHANGE Miyasaka et al. (2008) none 53.4 87 12.6
DE007 van de Putte et al. (2003) none 50.2 70 10.0
DE011 van de Putte et al. (2004) none 53.5 110 18.2
@@
Figure 3 illustrates how SPx would adjudicate between its three borrowing mechanisms over a continuum of possible new trial data, with the observed response rate ranging from (roughly) 10% to 50%. Like in Figure 2, the new trial’s control arm has had previous MTX treatment and has an average age of 53. In panels A and B we can see that the degree of borrowing is largely controlled by how close the new trial’s rate is to the observed overall and MTX-specific historical rates. The $hist$ and $reg$ submodels receive their maximum posterior weighting near these rates. In contrast, as the observed rate diverges from these historical rates the no-borrowing submodel quickly gains posterior weight, especially when the new trial’s control group size is larger. Panel C shows how the SPx prior leads to different Stage 2 sample size determinations depending on the new trial’s observed response rate at interim. When the observed rate is close to the predictions of the $hist$ and $reg$ submodels, the Stage 2 sample size can be very small (compared to the target size of 75 with no borrowing). However, when the observed rate is far from these historical predictions (i.e. $<0.22$ or $>0.4$ on the x-axis), the design can require a slightly larger Stage 2 control group. In this case the modeling assumption to allow historical borrowing actually increases uncertainty about the new trial’s rate compared to not borrowing at all, so the design calls for collecting more data than otherwise planned. The profile of sample size decisions plotted in Panel C is influenced by the SPx model, and in particular the prior probability of the no-borrowing submodel, $p_{ind}$. If this profile were deemed unreasonable in practice, this may be a sign that the trialists should consider changing $p_{ind}$ to better reflect their beliefs.

### v2

@@ 3 Adaptive Design Based on Posterior Inference
We propose a two-stage adaptive design, largely following Schmidli et al. [2014], to reduce the control group size to the extent that reliable information can be gained through historical borrowing. Intuitively, in Stage 1 the new trial enrolls a fixed and prespecified number of control group patients. Then there is an interim check to determine the number of control patients to enroll in Stage 2. The interim check measures the degree of “compatibility” between the historical and new trial (Stage 1) data in terms of an effective sample size of control patients gained by borrowing from the historical data. If the new trial’s stage one control data is deemed less compatible with the historical data, little borrowing will happen and the Stage 2 size will not be reduced much or at all. This limits the impact of prior assumptions about the relevance of the historical data on the overall trial size. Because we model the treatment group data independently from the control group data (having no shared parameters), the treatment group patients may be enrolled without reference to the control patients or the stages of the adaptive design (i.e. by simply changing the randomization probability after the interim check).
To define the adaptive design, we introduce notation for the data from various sources and stages of the trial. Let the historical data be denoted as $D_{hist}=(y_{h},\bm{x}_{h},n_{h})_{h=1,\dots,H}$, let the new Stage 1 data (control and treatment) be denoted as $D_{1}=\left(y_{H+1,1},\phantom{,}\bm{x}_{H+1,1},\phantom{,}n_{H+1,1}^{c}\right)\phantom{,}\cup\phantom{,}\left(y_{trt,1},\phantom{,}\bm{x}_{trt,1},\phantom{,}n_{trt,1}^{c}\right)$, and let the new Stage 2 data (control and treatment, not including patients from Stage 1) be denoted as $D_{2}=\left(y_{H+1,2},\phantom{,}\bm{x}_{H+1,2},\phantom{,}n_{H+1,2}^{c}\right)\phantom{,}\cup\phantom{,}\left(y_{trt,2},\phantom{,}\bm{x}_{trt,2},\phantom{,}n_{trt,2}^{c}\right)$. The data we analyze at interim are $D_{I}=D_{hist}\cup D_{1}$ and the data we analyze at the final analysis are $D_{F}=D_{hist}\cup D_{1}\cup D_{2}$.
@@
    Collect data $D_{1}$ on $n_{trt,1}^{c}$ treatment patients and $n^{c}_{H+1,1}$ control patients.
  • Interim analysis:
    • (a)
      Compute the interim SPx posterior of $\psi_{H+1}$ given the interim data $D_{I}=D_{hist}\cup D_{1}$, i.e. $p\left(\psi_{H+1}|D_{I}\right)$.
    • (b)
      Find the effective sample size of this posterior, $n^{c}_{H+1,\phantom{,}eff}$, via the moment matching method of Weber (2020).
@@
    Collect data $D_{2}$ on $n_{trt,2}^{c}$ treatment patients and $n^{c}_{H+1,2}$ control patients, where $n^{c}_{H+1,2}=n^{c}_{H+1,max}-n^{c}_{H+1,eff}$ but is truncated to the interval $[\gamma_{min}n^{c}_{H+1,\phantom{,}max},\phantom{,}\gamma_{max}n^{c}_{H+1,\phantom{,}max}]$ for some $\gamma_{min}\in[0,1]$, $\gamma_{max}\in[1,\infty)$ (e.g. $\gamma_{min}=0.75$, $\gamma_{max}=1.25$).
  • Final analysis:
    Inference for the treatment effect $\delta=\psi_{trt}-\psi_{H+1}$ is based on the SPx posterior given the final data $D_{F}=D_{hist}\cup D_{1}\cup D_{2}$, i.e. $p\left(\delta|D_{F}\right)$.
The second stage is designed so that the total sample size of the control group is not intolerably lower or higher than the target size $n^{c}_{H+1,max}$.
We emphasize that inferences for $\psi_{H+1}$ and $\delta$ (as well as all other parameters) follow standard Bayesian principles at both interim and final analysis. In particular, the posterior distribution used for the interim analysis is
[…]
where $\bm{\psi}=\left(\psi_{1},\dots,\psi_{H},\phantom{,}\psi_{H+1},\phantom{,}\psi_{trt}\right)$ and $\bm{z}=(m_{hist},\phantom{,}m_{reg},\phantom{,}m_{ind})$ and $d\psi_{H+1}^{-}$ is shorthand to indicate that we integrate over all parameters in SPx other than $\psi_{H+1}$. The posterior of $\psi_{H+1}$ at the final analysis is defined similarly, with $p(D_{F}|\psi_{H+1})$ replacing $p(D_{I}|\psi_{H+1})$:
[…]
The third line in (6) expresses the final posterior as a standard Bayesian update of the interim analysis [Berry et al. 2010], where the new data is $D_{2}$ and the “prior” is the interim posterior, $p\left(\psi_{H+1}|D_{I}\right)$. The posterior for $\delta$ at the final analysis is induced by the posterior for $\psi_{H+1}$ and the independent posterior for $\psi_{trt}$ (as discussed in Section 2).
Crucial to this design is the posterior effective sample size of SPx at interim, $n^{c}_{H+1,\phantom{,}eff}$, which is intended to assess how much information the historical data contribute about the new trial’s control rate and how many fewer patients the new control group needs in exchange for this additional information [Hobbs et al. 2013, Schmidli et al. 2014]. We use a simple moment matching definition of effective sample size (Weber 2020), which finds the Beta distribution with the same mean and variance as the SPx posterior, takes the effective sample size of this Beta distribution (the sum of its parameters; Morita et al. [2008]), and subtracts the current sample size at interim, $n^{c}_{H+1,1}$. We note that it is technically possible that $n^{c}_{H+1,eff}<n^{c}_{H+1,1}$, indicating that using the historical data reduce certainty about the new trial’s control rate and the Stage 2 sample size should be slightly increased to compensate. In truth, defining an effective sample size that has desirable properties for complex non-conjugate models such as SPx is an open area of research [Morita et al. 2008; 2012, Neuenschwander et al. 2020]. However, the sample size thresholds used in Stage 2 of the design somewhat limit the impact of the choice of definition. The simple definition we use has the benefit of producing more conservative (smaller) effective sample sizes for SPx than other definitions, which sometimes produce implausibly large effective sample sizes for the SPx model. Regardless of the specific definition used, by measuring the effective sample size during, and not before, the trial we can measure the extent to which the new data diverge from the historical data, potentially safeguarding against inappropriate borrowing from irrelevant historical data.
@@
$P(\delta>0|D_{F})$ $\geq 1-q_{positive}$ (7)
for the treatment effect $\delta:=\psi_{trt}-\psi_{H+1}$ where $q_{positive}\in(0,1)$ is the posterior Type I error threshold. Alternatively, to detect clinically significant effects, we may use the rule
$P(\delta>\delta_{0}|D_{F})$ $\geq 1-q_{clinical}$ (8)
@@ 4 Simulation Study
We compared several methods including SPx, RMAP [Schmidli et al. 2014], and an independent model (Ind.) with no historical borrowing (i.e. $\theta_{H+1}\sim Logistic(0,1)$). For SPx and RMAP we included both versions where the new trial’s control group size was fixed and where the two-stage adaptive design described in Section 3 was used (with $n^{c}_{H+1,1}=n^{c}_{H+1,max}/2$, $\gamma_{max}=1.25$, and $\gamma_{min}=0.75$). The two versions gave a contrast on the potential gain in the adaptive design in reducing the control group sample size. The value $n^{c}_{max}$ is fixed at 200 or 100, which are the maximum control sample size. For SPx our standard implementation used only 2 of 6 covariates associated with response rates to mimic imperfect knowledge or data collection, but we also include cases where SPx uses all 6 covariates to show the effectiveness of the method when the regression is strong and not misspecified. For RMAP we used a 50-50 prior mixture between the borrowing (MAP) component and the non-borrowing component, since this specification performed well in Schmidli et al. [2014]. The 90-10 mixture they consider was too aggressive and extremely biased in many of our scenarios. We emphasize that like SPx, RMAP adaptively updates the posterior weights (away from the 50-50 weights in the prior) for the two components depending on their relative model fits. This is consequence of the standard BMA formulation, and can be seen in equations (7) and (8) of Schmidli et al. [2014] (similar to our equation (4)). For each method we computed the posterior mean and 95% credible interval for $\psi_{H+1}$, as well as the posterior probabilities of $\delta$ described in equations (7) and (8). Note that these estimates and tests are Bayesian, so the credible intervals are not expected to have exactly 95% Frequentist coverage and the tests are not expected to have exactly 5% type I error rates.
@@ 4.1 Estimation Accuracy for the New Trial’s Control Rate
RMSE 0.018 0.018 0.026 0.028 0.029 0.022 0.022 0.031 0.033 0.039
Coverage 98.8 98.9 95.0 96.4 93.9 99.3 99.2 97.6 97.8 95.8
Width 0.086 0.094 0.103 0.112 0.110 0.117 0.130 0.138 0.152 0.155
@@
Table 1: Control group size and Frequentist estimation accuracy for the new trial’s control response rate, averaged over 1,000 simulated trials. Metrics are defined as follows: size is the mean control group size; RMSE is the root mean square error of the posterior mean of the control group rate; coverage is the proportion of trials in which the 95% quantile-based posterior credible interval for the control group rate contains the true rate; width is the mean width of these credible intervals. Note that the credible intervals are Bayesian and not designed or calibrated to give exactly 95% Frequentist coverage. For the Ind Fixed method, the theoretical operating characteristics are identical in (a) Scenarios 1 and 3 (for a given sample size) and in (b) Scenarios 2 and 4 (for a given sample size) so we averaged results in these pairs.
@@ 4.2 Power and Type I Error for the New Trial’s Treatment Effect
Table 2 examines the statistical power of the adaptive designs assuming 2 and 6 covariates were included in SPx, respectively. Type I error is relatively well controlled by all methods. Of course, a drug developer or regulator requiring that Frequentist Type I error be more strictly controlled can calibrate the model or decision rule by simulation under the specific scenarios they are concerned about. This is a reality of all Bayesian trial methods and many Frequentist ones as well [e.g. Lewis et al. 2007], a point we revisit in the Discussion.
@@
0 $P(\delta>0)$ 5.5 5.2 6.6 5.9 4.9 6.2 6.3 8.2 8.8 5.7
0.3 $P(\delta>0.2)$ 69.8 70.2 70.1 71.4 69.1 45.2 45.7 47.8 50.1 43.1
Table 2: Frequentist Type I error and power for the new trial’s treatment effect, averaged over 1,000 simulated trials. For convenience, the mean control group size is repeated from Table 1. Type I error is shown in the $P(\delta>0)$ rows, where each model declares a positive treatment effect when its posterior probability that $\delta>0$ exceeds 95%; this row shows the % of simulated trials where $\delta=0$ in which each method instead declares $\delta>0$. Power is shown in the $P(\delta>0.2)$ rows, where each model declares a clinically significant treatment effect when its posterior probability that $\delta>0.2$ exceeds 95%; this row shows the % of simulated trials where $\delta=0.3$ in which each method declares $\delta>0.2$. For the Ind Fixed method, the theoretical operating characteristics are identical in (a) Scenarios 1 and 3 (for a given sample size) and in (b) Scenarios 2 and 4 (for a given sample size) so we averaged results in these pairs.
@@ 4.3 SPx’s Adaptive Weighting of Borrowing Mechanisms
To illustrate how SPx automatically adjusts the type of borrowing it performs depending on the historical and new trial data, Figure 2 plots the simulation distribution of SPx’s posterior submodel weights $p^{*}_{k}$ (see equation (3)) in Scenarios 1 through 4. When the historical data are entirely misleading (Scenario 4) SPx strongly favors its no-borrowing submodel. Otherwise SPx puts most posterior mass on its two borrowing submodels, appropriately favoring the regression submodel over the historical one when the historical rates are misleading but covariates are useful (Scenario 2).
In the reversed setting where the historical rates are not misleading but covariates are not predictive (Scenario 3) SPx still gives moderate weight to the regression model because its inclusion of covariates adds some noise but not substantial bias. In particular, in Scenario 3 SPx correctly identifies that covariates are not predictive, meaning that the posterior distribution for the coefficient vector $\beta=(\beta_{0},\beta_{1},\dots,\beta_{p})$ concentrates on a vector $(b_{0},0,\dots,0)$ for some $b_{0}\in\mathbb{R}$. The first element in this vector is the intercept in the regression function, representing the overall mean of the (logit) control response rates across studies. The remaining elements are the coefficients for the covariates, all of which are not associated with the outcome. In other words, in Scenario 3 the posterior of the regression submodel is similar to the simple MAP model shown in Section 2.1. This provides a similar inference on $\theta_{H+1}$ to the historical submodel, which also assumes that $\theta_{H+1}$ is close to a measure of the central tendency of $\theta_{1},\dots,\theta_{H}$ (i.e. their weighted sample mean). Thus, in Scenario 3, it is sensible and expected that the regression and historical submodels would receive comparable weight.
@@
It is also important to note that even when the historical data are relevant (Scenarios 1-3), in our simulations the independent no borrowing model still received non-trivial posterior weight (on average  20% - 50% depending on the scenario). This is because SPx puts a conservative prior on the submodel weights, with 75% prior probability going to the no-borrowing model. However, because the direct historical and regression submodels make aggressive predictions that heavily trust the historical data (see Section 3 and Supplementary Section S3), SPx still borrows significantly from the historical data in these scenarios. This is reflected in the increased efficiency of SPx’s operating characteristics for Scenarios 1-3 in Tables 1 and 2. In this sense, the posterior submodel weights should not be interpreted by themselves as the sole summary of SPx’s borrowing behavior.
4.4 SPx’s behavior when regression is biased for the new trial
To illustrate how the regression and direct historical submodels contribute differently to SPx’s Bayesian model averaging, we conducted additional simulations in which the regression submodel is correctly specified for the historical trial data but provides biased predictions for the new trial. These scenarios reflect situations where, for example, unmeasured confounding or systematic differences in covariate measurement make the historical regression function invalid in the new trial.
Scenarios 5 and 6 are modifications of Scenario 2. In Scenario 2, the regression model is correctly specified for both the historical and new trials. However, the new trial has extreme covariate values, so its true response rate is 44.5% compared to an average of 23% in the historical trials. In the new scenarios, we manually changed the new trial’s true response rate while keeping the regression model, the historical trial data, and the new trial’s covariates fixed. In Scenario 5, the new trial’s response rate is 26%, which is close to the historical average. In Scenario 6, it is 33%, which is farther from the historical average but still well below the regression prediction of 44.5%. These changes make the regression model correctly specified for the historical trials but misspecified for the new trial.
The results for Scenarios 5 and 6 (Supplementary Figure S2) show that SPx correctly and substantially down-weights the regression submodel, relying more heavily on the direct historical submodel (and, to a lesser extent, the independent submodel, especially in Scenario 6). As a result, SPx continues to provide accurate estimates of the new trial’s control response rate (Supplementary Table S2) and accurate tests of its treatment effect (Supplementary Table S3). To summarize, in these scenarios the regression submodel is badly biased for the new trial. However, SPx limits its influence through Bayesian model averaging. SPx automatically identifies that the regression submodel is less predictive of the new trial’s observed data (in terms of marginal likelihood, see equation (4)) and thus shifts its weight to better-performing alternatives, preserving the accuracy of the reported BMA inferences.
We note that for readers concerned about functional form assumptions (i.e. linearity) in the regression submodel, these simulation results offer some reassurance. What matters in SPx is not whether the regression submodel is strictly correct, but whether it provides accurate predictions compared to the other submodels, as measured by the marginal likelihoods. If a bad functional form assumption makes the regression submodel sufficiently inaccurate, it will be down-weighted in the SPx posterior in favor of better-performing submodels.
In settings where the new trial’s covariates are clear outliers relative to the historical controls’ covariates, extrapolation is unavoidable. In such cases, analysts should carefully consider whether the historical data should be used at all, regardless of the borrowing method. Even methods that do not explicitly model the response-covariate relationship can suffer from inappropriate extrapolation. For example, in propensity score methods, outlier covariate values in the new trial can violate the positivity (or common support) assumption [Westreich and Cole 2010]. This substantially complicates the analysis, and the standard guidance is to remove units with extreme propensity scores — effectively recommending that the historical and new trial data not be combined [Petersen et al. 2012, Hill and Su 2013, Zhou et al. 2020]. SPx provides some automatic protection against such extrapolations because it will revert primarily to the independent submodel (see Supplementary Figure S2). However, if large covariate discrepancies between the new and historical trials are expected in advance or are observed in the data, we advise against using historical data for design and analysis, regardless of the method used.
5 Rheumatoid Arthritis Case Study
@@
We discuss the application of SPx to the development of novel treatments for rheumatoid arthritis (RA). RA is an auto-immune disease that affects more than 1.3 million patients in the United States (Hunter et al. 2017), and it is predicted that its global burden will increase through 2030 [Cai et al. 2023]. The symptoms of this disease include pain and swelling joints in hands and feet as well as morning stiffness lasting longer than 30 minutes. Although there is no cure for RA, several treatments are available to slow down the disease progression and alleviate symptoms. Adalimumab is a well-established, standard biologic therapy in RA and has been approved for 20 years. As the patent of adalimumab is expiring globally and becoming a standard of care, novel therapies for RA must demonstrate superior treatment effects over adalimumab. We apply our SPx methodology in adalimumab trials conducted in the past two decades to explore possibilities for the design and analysis of a future trial in RA. Since adalimumab’s own development was extensive, there are many past trials including an adalimumab arm that could be used as a rich source of historical data to potentially accelerate trials of novel RA drugs. In this process, SPx uses trial-level covariate information and a combination of historical borrowing mechanisms to guide the new trial.
@@
ALTARA Kennedy et al. [2014] MTX 48.8 43 39.5
ARMADA Weinblatt et al. [2003] MTX 56.0 62 21.0
DE019 Keystone et al. [2004] MTX 56.1 200 24.0
IM133-001 Weinblatt et al. [2015] MTX 51.4 61 39.3
ORAL-Standard van Vollenhoven et al. [2012] MTX 53.7 106 26.4
RA-BEAM Taylor et al. [2017] MTX 53.0 488 40.2
STAR Furst et al. [2003] MTX 55.8 315 29.5
A3921035 Fleischmann et al. [2012] none 53.0 59 22.0
CHANGE Miyasaka and The CHANGE Study Investigators [2008] none 53.4 87 12.6
DE007 van de Putte [2003] none 50.2 70 10.0
DE011 van de Putte [2004] none 53.5 110 18.2
@@
Figure 4 illustrates how SPx would adjudicate between its three borrowing mechanisms over a continuum of possible new trial data, with the observed response rate ranging from (roughly) 10% to 50%. Like in Figure 3, the new trial’s control arm has had previous MTX treatment and has an average age of 53. In panels A and B we can see that the degree of borrowing is largely controlled by how close the new trial’s rate is to the observed overall and MTX-specific historical rates. The $hist$ and $reg$ submodels receive their maximum posterior weighting near these rates. In contrast, as the observed rate diverges from these historical rates the no-borrowing submodel quickly gains posterior weight, especially when the new trial’s control group size is larger. Panel C shows how the SPx prior leads to different Stage 2 sample size determinations (based on the adaptive design described in Section 3) depending on the new trial’s observed response rate at interim. When the observed rate is close to the predictions of the $hist$ and $reg$ submodels, the Stage 2 sample size can be very small (compared to the target size of 75 with no borrowing). However, when the observed rate is far from these historical predictions (i.e. <0.22<0.22< 0.22 or >0.4>0.4> 0.4 on the x-axis), the design can require a slightly larger Stage 2 control group. In this case the modeling assumption to allow historical borrowing actually increases uncertainty about the new trial’s rate compared to not borrowing at all, so the design calls for collecting more data than otherwise planned. The profile of sample size decisions plotted in Panel C is influenced by the SPx model, and in particular the prior probability of the no-borrowing submodel, $p_{ind}$. If this profile were deemed unreasonable in practice, this may be a sign that the trialists should consider changing $p_{ind}$ to better reflect their beliefs.